OpenStudy (anonymous):

Had the answer wrote it down and lost its. find the slope if one exists 5x + 6Y = 30. Please tel me how you came up with the answer????

OpenStudy (amistre64):

what does slope measure?

OpenStudy (amistre64):

i mean, there is nothing special about this problem that makes it defy all known slope related conundrums :)

OpenStudy (anonymous):

-5/6 i think

OpenStudy (amistre64):

6/30 maybe ... but how does throwing out an unsubstatiated number "help"?

OpenStudy (anonymous):

The numbers you all see is what I have and I am in the dark on how to do it

OpenStudy (anonymous):

you want to isolate the y on one side of the equation

OpenStudy (anonymous):

so 6y=-5y+30

OpenStudy (amistre64):

isolating they is good; if you know where to find the slope at in it :)

OpenStudy (anonymous):

then divide both sides my 6 so y=(-5/6)y+30

OpenStudy (anonymous):

because you put the equation into the form y=mx+c (the formula of a straight line) and that gives you 6x=30-5x Then divide everything by 6 and there ya go:)

OpenStudy (anonymous):

the form is y=mx+b, b being the slope so the slope is (-5/6)

OpenStudy (anonymous):

sorry m is slope

OpenStudy (amistre64):

but the slope can be found by the formula: ax+by=c just as well right? slope = -a/b in that case

OpenStudy (amistre64):

but knowing what slope measures is important. When you know what it measures, finding it is a breeze

OpenStudy (amistre64):

slope is: the change in "y" with respect to a change in "x" 5x + 6Y = 30 ; lets make x=0 and solve for y 5(0) + 6y = 30 6y = 30 y = 30/6 = 5 now lets change "x" by 1; say ...x=1 5(1) + 6Y = 30 5 + 6y = 30 6y = 30 - 5 6y = 25 y = 25/6 now we see how y moved and that is our slope: y went from 5 to 25/6; or rather: from 30/6 to 25/6; which is a movement of -5/6 our slope is -5/6

OpenStudy (anonymous):

If this is an algebra question then... Place the equation in standard form (y=mx+b) then the coefficient (number) in front f the x variable is the slope. 5x+6y=30 is not in standard form. To get a linear equation into standard form: 1) Move the x variable and its coefficent to the rigth side of the equation using an arithmetic operation (+,-). 2) Get y all by itself on the left side with a coefficient of 1 by using another arithmetic operation (x./). example: 2y +4x =10 2y+4x-4x=10-4x :subtraxt 4x from both sides 2y=10-4x 2y/2=(10-4x)/2 :divide both sides by 2 y=5-2x the coefficient in front of the x variable is -2. Therfore the slope is -2